jones can plough his field with his tractor in 4 hours. if he uses his manual plough, it takes 3 times as long to plough the same field. one day after working with the tractor for 2 hours, he ran out of gas and had to finish with the manual plough. how long did it take to complete this job after the tractor ran out of gas
Answer by richwmiller(17219)

He finished half of the job in those two hours 2/4=1/2
6/12 =1/2 so he has 6 more hours to go instead of 2 with the tractor.
or
3*2=6 more hours to go